f(x) = √(cotx.x³) dx
1
Expert's answer
2020-05-31T15:14:08-0400
"f(x)=\sqrt{\cot x\cdot x^3}"

"{df \over dx}={1 \over 2\sqrt{\cot x\cdot x^3}}(-{x^3 \over \sin^2x}+3x^2\cot x)="

"={\sqrt{x}(3\sin x\cos x-x) \over 2\sqrt{\cos x\sin ^3x}}"